OpenStudy (anonymous):

Simplify the rational expression. Assume that the denominator is not 0 32p^5q^2 over 12p^6q

OpenStudy (anonymous):

(32p^5q^2)/(12p^6q)=32*p^5*q^2*(12^-1)*(p^-6)*q^(-1)=8q/p

OpenStudy (anonymous):

answer is 8q/p?

OpenStudy (anonymous):

its wrong!